1. (15 pts) Evaluate the following integral: 171*** v2 + y2 +7 derdy de
Answer #1

3 19-02 19-x²y2 Sob Jabytez de dy dre ue spherical coordenak: ocne3 27 os y = 19-22 3 y = 19-42 t Hemisphere x + y + z = 82/